Today Mark got ahold of and installed a new diode into one of the -5V rails of Chamber 4 right before the sweep.
A replay of a cosmic run taken right before the sweep showed two gaps in the number of hits v wire group histogram for U9, U10, U11 wire groups 9-10 and 19-24. After the sweep we went into the Hall to check out what could be causing this. It was during this investigation that we noticed that Pwr Supply 13c read +3V when it should be -5V (although using a DVM, we measured -5V).
We also noticed that the modules in which these "missing" wire groups lie did not seem to have an amplified signal (the signal was just a pulse, not multiplexed). This would seem to indicate that the cards in the chamber are not receiving the appropiate power and are therefore not amplified and therefore not seen in the data.
Mark Jones has been notified and we will attack this problem tomorrow.
FPP Chamber 4 from online GUIFPPOnlineGUI_Chamber4_11Feb2011.eps